If sinx+siny=a and cosx+cosy=2a what is the value of cos(x-y) ?

Grade:12

1 Answers

Bhumiti
7 years ago
We know that cos(x-y) = cosx.cosy+sinx.siny... So just substitute the given values in the above equation... We get cos(x-y)=2a+a =3a